# To Offload or Not To Offload: Model-driven Comparison of Edge-native and On-device Processing

### v1

@@ Abstract.
submissionid: 696copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXX
1. Introduction
@@
Over the past decade, cloud computing has become the predominant approach for running online services in domains ranging from finance to entertainment. In recent years, a new class of online services has emerged that requires low-latency processing to meet end-user requirements. Examples of such applications include autonomous vehicles, interactive augmented and virtual reality (AR/VR), and human-in-the-loop machine learning inference (Mosqueira-Rey et al., 2022). Edge computing, which is a complementary approach to cloud computing, is a promising method for meeting the needs of such latency-sensitive applications. Conventional wisdom has held that since edge resources are deployed at the edge of the network and are closer to end-users, edge-native processing can provide lower end-to-end latencies than cloud processing. However, recent research (Ali-Eldin et al., 2021; Wang et al., 2024a) has shown that this conventional wisdom does not always hold—in certain scenarios, edge processing can yield worse end-to-end latency despite its network latency advantage over the cloud. In particular, since edge clusters are resource-constrained relative to the cloud, research has shown resource bottlenecks can arise in the presence of time-varying workloads and increase the processing latency at the edge, negating its network latency advantage. Further, such bottlenecks can arise even at moderate levels of utilization, requiring careful resource management to maintain the benefits of edge-native processing (Wang et al., 2024a).
@@
A related trend is the emergence of computational offloading from local devices to remote resources such as those at the edge. [^1] Computational offloading to the cloud is also common, but we focus on edge offloading in this paper. Computational offloading (Satyanarayanan et al., 2009; Davis et al., 2004) involves using nearby edge resources to take on some, or all, of an application’s processing demands. Such offloading is performed to overcome resource constraints on a local device by leveraging more abundant processing capacity at edge servers. Computational offloading is well-established in mobile computing where mobile devices, which are resource- or battery-constrained, offload local computations to edge servers (Ren et al., 2019). In this case, local devices pay a so-called mobility penalty (Satyanarayanan et al., 2019), which is the additional network latency to access edge servers, but then benefit from faster edge processing that outweighs this latency overhead. Offloading is also common for Internet-of-Things (IoT) devices that have limited on-board processing abilities and rely on remote processing to implement data processing tasks (Jang et al., 2018). In these offloading scenarios, conventional wisdom holds that remote edge processing is faster than local processing despite the network overhead of accessing remote resources due to significant resource constraints at local devices.
@@
With the advent of hardware accelerators in recent years, the conventional wisdom about the benefits of computational edge offloading for local devices needs to be rethought. For example, the rise of the so-called AI PC with onboard AI accelerators (Intel Corporation, 2025) has enabled local AI processing for many applications without a need for edge offloading. Similarly, mobile phones are equipped with neural accelerators (Simonite, 2017) that enable efficient local inference, while IoT devices such as cameras have specialized accelerators to perform visual tasks (e.g., facial recognition) locally (Netatmo, 2025). At the same time, edge servers can also be equipped with more powerful accelerators, such as GPUs, enabling offloaded tasks to also run more efficiently or faster than before. Thus, in the era of programmable accelerators, the question of whether to offload computations from local devices to the edge and when it is advantageous to do so needs to be reexamined. In particular, such a reexamination needs to address several questions. (i) Under what scenarios will local on-device processing using accelerators outperform edge offloading? (ii) Are there still scenarios where edge offloading has an advantage over local processing? (iii) What role does a fast (or slow) network between the device and the edge play in terms of whether to offload or not offload?
@@
Motivated by the above questions, this paper takes an analytic model-driven approach to compare local and edge processing in the era of accelerators. Our hypothesis is that there is no a priori clear advantage of edge processing over on-device processing in the presence of accelerators, and the choice of where to perform the processing will depend on hardware, network, and workload characteristics. In developing our analytic models and validating this hypothesis, this paper makes the following contributions.
@@
  • We use queuing theory-based models to model the behavior of device and edge accelerators when processing requests for local on-device processing and edge offloading scenarios. Using these models, we develop closed-form analytic bounds to determine when one approach outperforms the other under different hardware, network, and workload scenarios. We also develop analytic models for comparing on-device processing to multi-tenant edge servers that service multiple clients.
@@
  • We experimentally validate our analytic models and bounds using several different device and edge accelerators and network configurations and show that our models are able to predict the performance behavior of on-device processing and edge offloading under various scenarios. Our evaluation shows that our model achieves a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5of the observed latency and 100% within $\pm$10%.
@@
  • We develop a dynamic resource manager that incorporates our models and demonstrate the ability of this resource manager to adaptively switch between on-device and edge processing under time-varying workloads and under dynamic multi-tenancy scenarios. Our results show that it effectively adapts to changing network conditions and fluctuating request rates at edge servers.
2. Background
@@
Table 1. Characteristics of common programmable accelerators.
@@
2.1. Cloud and Edge Offloading
Computational offloading is a well-known approach where client devices that are resource-constrained offload some or all of their application workload to a remote server. In the case of mobile devices such as AR/VR headsets, tablets, or smartphones, offloading can use nearby edge resources to perform latency-sensitive tasks (Satyanarayanan et al., 2009; Davis et al., 2004). While such edge offloading involves a network hop to the edge server (also known as the mobility penalty (Satyanarayanan et al., 2019)), subsequent processing at the edge is assumed to be much faster since edge resources are much less constrained than the devices. In the case of IoT devices or ML applications, offloading to the cloud is also common where the abundant processing capacity of the cloud platform is utilized for processing the application workload (Huang et al., 2020). For example, large language model (LLM) inference uses GPU clusters in the cloud to answer user queries, while IoT devices use cloud storage and processing to perform analytic and data processing tasks.
@@ 2.1. Cloud and Edge Offloading
2.2. Hardware Accelerators
In recent years, a wide range of programmable accelerators has been developed to accelerate a range of compute tasks, particularly machine learning compute. Such accelerators have been deployed on client devices, edge, and cloud servers, enabling faster processing of accelerated tasks at each tier. Table 1 depicts the characteristics of several commonly deployed device and edge accelerators, ranging from low-end to high-end options. Specifically, the Intel Movidius Myriad X Vision Processing Unit (VPU) is designed to accelerate vision-based applications and CNN inference (Intel, 2017). Google Edge Tensor Processing Unit (TPU) optimizes ML inference with the TensorFlow Lite framework on the edge (Google, 2019). NVIDIA TX2, Orin Nano, and A2 are all tailored for AI workloads such as deep learning and computer vision (NVIDIA, 2017, 2023, 2021). For these specific workloads, programmable accelerators can provide substantial performance improvements, delivering server-like capabilities for some compute-intensive tasks even on low-end devices. With their small form factor and energy-efficient design, integrating accelerators into mobile devices has become increasingly common. For example, Apple devices include a Neural accelerator to accelerate machine learning tasks locally on the device (Simonite, 2017). With the presence of accelerators, fast on-device processing of tasks such as machine learning inference has become feasible without the need to resort to edge offloading. At the same time, edge servers can also be equipped with GPUs, which serve as general-purpose accelerators for a range of tasks, or special-purpose accelerators for specific tasks. This enables offloaded tasks to also be accelerated at the edge, making edge offloading useful for more compute-intensive tasks that a device may not be able to handle using local resources.
2.3. On-device Processing versus Edge Offloading
@@
Even with extensive offline profiling of execution times on both the device and the edge, selecting the optimal execution strategy remains challenging due to dynamic real-world conditions. On the one hand, mobile devices are subject to network variability due to fluctuating wireless signal strength, interference, and congestion in shared communication channels. Since edge offloading requires transmitting input data to an edge server, network fluctuations introduce unpredictability in network delays, affecting the overall offloading performance. Additionally, mobile devices running applications with prolonged operating durations, such as live object detection on captured video, may experience battery constraints. When operating in low-power mode, the system may underclock its processing units to conserve energy, resulting in extended execution times. On the other hand, edge servers are multi-tenant systems shared by multiple devices, which can result in interference across users and higher latencies. Since devices may independently decide when and whether to offload tasks, the edge server’s load can fluctuate dynamically based on the number and nature of tasks being offloaded at any given time. Compared to traditional cloud data centers, edge servers are often more resource-constrained, limiting their ability to elastically scale applications under resource pressure. Workload spikes in multi-tenant scenarios can increase queuing delays, significantly increasing end-to-end execution times. Moreover, edge servers are commonly multiplexed across different applications to maximize edge resources utilization. At the same time, GPUs and other programmable accelerators designed for edge servers often lack isolation mechanisms such as GPU virtualization that are available in cloud-based GPUs (NVIDIA, 2025), leading to performance interference between applications. As a result, execution times can become variable, making it challenging to determine whether to offload under dynamic workloads.
@@
3. Model-driven Performance Comparison
In this section, we develop analytical queuing models for edge offloading and on-device processing and then develop closed-form expressions for their expected latencies to compare the two approaches and determine when one strategy outperforms the other.
3.1. Analytic Queuing Models
(a) Edge offloading queueing model
(b) On-device processing queueing model
Figure 1. Modeling request execution using (a) edge offloading and (b) on-device processing.
Our analytical models are derived from queuing theory, a well-established mathematical tool for modeling response times experienced by requests in computing systems. Queuing theory has previously been used to model the performance of edge and cloud applications (Ali-Eldin et al., 2021; Harchol-Balter, 2013; Urgaonkar et al., 2005), and will serve as the foundation for analyzing edge offloading and on-device processing. As discussed below, we model the end-to-end latency of a request—the duration of time from when the task enters the first queue until it exits the system—for both edge offloading and on-device processing using queuing theory.
Edge offloading. To model the end-to-end request execution time using edge offloading, we need to model the times spent by the request both at the device and at the edge server. We do so by modeling the device and the edge as two separate queuing systems as shown in Figure 1(a). We assume that the request arrives at the device and is then forwarded (“offloaded”) to the edge server over the network. We need to model any network queuing delays and the network transmission time of the request from the device to the server. Upon arrival at the edge server, the request is queued for processing and is subsequently scheduled for execution. Since accelerators employ data parallelism to process requests, we model the accelerator at the edge server as a parallel system that processes each request with a degree of parallelism $k_{edge}$. Once the request processing completes, the result is sent back to the device, which incurs network queuing delay and transmission latency on the reverse path.
On-device processing. Similar to edge offloading, we model on-device request processing as a queuing system shown in Figure 1(b). Incoming requests generated by the device first enter a local device queue and are then scheduled onto one of the accelerator cores of the device where they are processed with a degree of parallelism $k_{dev}$. Upon completion of processing, the result is returned to the local application and the request exits the system. This model captures an application where a client device generates tasks, utilizes its accelerator cores for local execution, and sends the result to the application.
@@ 3.1. Analytic Queuing Models
With the above notations, we now model the end-to-end latency of requests executed using each strategy. First, we have
(1) $T_{edge}=w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$
@@
(2) $T_{dev}=w_{dev}^{proc}+s_{dev}$
where $T_{edge}$ and $T_{dev}$ denote the end-to-end latency of requests executed using edge offloading and on-device processing, respectively.
3.2. Deriving and Analyzing End-to-End Latency
In this section we compute the expected latency for edge offloading and on-device processing. Our analysis focuses on applications that require processing results on the device, such as AR/VR applications. Note that the proposed analytic models can also capture the performance of applications requiring results on the edge, and our findings can be extended to these cases. In the following, we leverage classical queuing theory to fundamentally understand the conditions under which one strategy outperforms the other.
3.2.1. Comparing Edge Offload and On-Device Execution
We begin our analysis by examining the conditions under which edge offloading results in higher latency compared to on-device processing. For workloads that benefit from hardware acceleration, such as those utilizing GPUs or TPUs, processing times can be significantly reduced compared to processing on general-purpose CPUs. In the following discussion, we refer to such workloads as accelerator-driven workloads. Table 2 summarizes the notation used in our models.
Table 2. Notations used in analytic models.
@@ 3.2.1. Comparing Edge Offload and On-Device Execution
Lemma 3.0.
For accelerator-driven workloads, edge offloading incurs a larger average end-to-end latency than on-device processing when {dmath} s_dev - s_edge ¡ λμdevnetdevnet- λ) + λμedgenetedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc )
@@
(3) $w_{dev}^{proc}+s_{dev}<w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$
@@
(4) $s_{dev}-s_{edge}<w_{dev}^{net}+w_{edge}^{net}+\frac{D_{req}}{B}+\frac{D_{res}}{B}+w_{edge}^{proc}-w_{dev}^{proc}$
As shown in (Liang et al., 2023), workloads on programmable accelerators typically exhibit deterministic service times (denoted as $D$), as the number of operations per request remains constant. Since typical accelerators have a small degree of parallelism $k$ (Amert et al., 2017), the queuing system can be modeled as an $M/D/k/FCFS$ system, where request arrivals follow a Poisson process. While no closed-form solution exists for the expected wait time in such a system, a common approximation is to aggregate the service rate across the degree of parallelism $k$ (Shah et al., 2018, 2020). This reduces the model to an $M/D/1/FCFS$ system with aggregated service rate $k\mu$. Its expected wait time can be derived from the well-known Pollaczek-Khinchin (P-K) formula (Gross et al., 2008; Harchol-Balter, 2013) and is given by
(5) $E[w_{M/D/1}]=\frac{1}{2}\left(\frac{1}{k\mu-\lambda}-\frac{1}{k\mu}\right)$
@@
(6) $E[w_{M/M/1}]=\frac{1}{\mu-\lambda}-\frac{1}{\mu}$
@@
Moreover, regardless of the execution strategy, the application sees the same arrival rate, i.e., $\lambda_{dev}^{proc}=\lambda_{edge}^{proc}=\lambda$. Substituting the above into (3.2.1), we obtain: {dmath} s_dev - s_edge ¡ λμdevnet(μdevnet- λ) + λμedgenet(μedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc ) which completes the proof. ∎
@@
Explanation. Service time, which is the time spent processing a request on an accelerator, is determined by the computational demand of the workload divided by the processing capacity of the hardware. For any given hardware configuration, requests with lighter demand have to be processed much faster at the edge to overcome the network penalty for edge offloading to be faster than on-device processing. For requests with lower computational demands, the difference in average service times $s_{dev}-s_{edge}$ shrinks proportionally due to the disparity in processing capacities between the device and the edge as the workload computational demand decreases. Further, since the service rate is the inverse of service times in queuing theory (i.e., $s=1/\mu$), the terms with $\mu_{dev}^{proc}$ and $\mu_{edge}^{proc}$ on the right hand side of the inequality in (3.1) will also become smaller since $\mu$ increases for lighter requests. The terms representing the network overheads remain unchanged since they depend on request size rather than the processing demand of the request. As a result, the inequality in (3.1) is more likely to hold true for requests with lighter processing demand. ∎
Remark 3.3.
On-device processing is more likely to outperform edge offloading for slower networks or when workloads have larger request or result payloads.
Explanation. The term $\frac{D_{req}+D_{res}}{B}$ in (3.1) increases when either the network bandwidth $B$ decreases or the request and result sizes $D_{req}$ and $D_{res}$ become larger, thereby increasing the right hand size of the inequality. In addition, longer transmission times reduce the service rates of the NIC at the device and edge ($\mu_{dev}^{net}$ and $\mu_{edge}^{net}$), further increasing the RHS. As a result, the inequality is more likely to hold for slower networks or tasks with higher network overheads, making on-device processing faster than edge offloading. ∎
Practical takeaways. Lemma 3.1 establishes a bound on the average service time difference between device and edge for edge offloading to be more effective. The bound and the resulting inequality depends on several factors: (i) relative processing speeds of the device and edge, denoted by $s_{dev}$, $s_{edge}$ (and also the corresponding service rates $\mu_{dev}$ and $\mu_{edge}$), (ii) network speed, (iii) overall workload denoted by request rate $\lambda$, (iv) network transmission overheads. In general, for any specific workload, the edge processing capacity must be larger than the device processing capacity by a factor that overcomes the network penalty of offloading the workload; otherwise, on-device processing becomes preferable to offloading.

### v2

@@ Abstract
publicationid: pubid:
I Introduction
@@
Over the past decade, cloud computing has become the predominant approach for running online services in domains ranging from finance to entertainment. In recent years, a new class of online services has emerged that requires low-latency processing to meet end-user requirements. Examples of such applications include autonomous vehicles, interactive augmented and virtual reality (AR/VR), and human-in-the-loop machine learning inference [1]. Edge computing, which is a complementary approach to cloud computing, is a promising method to meet the needs of such latency-sensitive applications. Conventional wisdom has held that since edge resources are deployed at the edge of the network and are closer to end-users, edge-native processing can provide lower end-to-end latencies than cloud processing. However, recent research [2, 3] has shown that this conventional wisdom does not always hold—in certain scenarios, edge processing can yield worse end-to-end latency despite its network latency advantage over the cloud. In particular, since edge clusters are resource-constrained relative to the cloud, research has shown resource bottlenecks can arise in the presence of time-varying workloads and increase the processing latency at the edge, negating its network latency advantage. Further, such bottlenecks can arise even at moderate levels of utilization, requiring careful resource management to maintain the benefits of edge-native processing [3].
@@
A related trend is the emergence of computational offloading from local devices to remote resources such as those at the edge. [^1] Computational offloading to the cloud is also common, but we focus on edge offloading in this paper. Computational offloading [4, 5] involves using nearby edge resources to take on some, or all, of an application’s processing demands. Such offloading is performed to overcome resource constraints on a local device by leveraging more abundant processing capacity at edge servers. Computational offloading is well-established in mobile computing where mobile devices, which are resource- or battery-constrained, offload local computations to edge servers [6]. In this case, local devices pay a so-called mobility penalty [7], which is the additional network latency to access edge servers, but then benefit from faster edge processing that outweighs this latency overhead. Offloading is also common for Internet-of-Things (IoT) devices that have limited on-board processing abilities and rely on remote processing to implement data processing tasks [8]. In these offloading scenarios, conventional wisdom holds that remote edge processing is faster than local processing despite the network overhead of accessing remote resources due to significant resource constraints at local devices.
@@
With the advent of hardware accelerators in recent years, the conventional wisdom about the benefits of computational edge offloading for local devices needs to be rethought. For example, the rise of the so-called AI PC with onboard AI accelerators [9] has enabled local AI processing for many applications without a need for edge offloading. Similarly, mobile phones are equipped with neural accelerators [10] that enable efficient local inference, while IoT devices such as cameras have specialized accelerators to perform visual tasks (e.g., facial recognition) locally [11]. At the same time, edge servers can also be equipped with more powerful accelerators, such as GPUs, enabling offloaded tasks to also run more efficiently or faster than before. Thus, in the era of programmable accelerators, the question of whether to offload computations from local devices to the edge and when it is advantageous to do so needs to be reexamined. In particular, such a reexamination needs to address several questions. (i) Under what scenarios will local on-device processing using accelerators outperform edge offloading? (ii) Are there still scenarios where edge offloading has an advantage over local processing? (iii) What role does a fast (or slow) network between the device and the edge play in terms of whether to offload or not offload?
@@
Motivated by the above questions, this paper takes an analytic model-driven approach to compare local and edge processing in the era of accelerators. Our hypothesis is that there is no a priori clear advantage of edge processing over on-device processing in the presence of accelerators, and that the choice of where to perform the processing depends on hardware, network, and workload characteristics. In developing our analytic models and validating this hypothesis, this paper makes the following contributions.
@@
  • We use queuing theory-based models to model the behavior of device and edge accelerators when processing requests for local on-device processing and edge offloading scenarios. Our modeling follows a two-level approach: it takes estimated service times obtained through profiling or other prediction techniques as input and embeds them into queuing models to obtain end-to-end response time predictions. Using these models, we develop closed-form analytic bounds to determine when one approach outperforms the other under different hardware, network, and workload scenarios. We also develop analytic models for comparing on-device processing to multi-tenant edge servers that serve multiple clients. Although our primary focus is on offloading, we also show how the models can be extended to account for device-edge collaborative (i.e., split) processing.
@@
  • We experimentally validate our analytic models and bounds using a variety of device and edge accelerators, network configurations, and workloads including deep neural networks (DNNs) and recurrent neural networks (RNNs). Our results show that our models are able to predict the performance behavior of on-device processing and edge offloading under various scenarios, achieving a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5of the observed latency and 100% within $\pm$10%.
@@
  • We develop a dynamic resource manager that incorporates our models and demonstrate how they can be leveraged by this resource manager to adaptively switch between on-device and edge processing under time-varying workloads and under dynamic multi-tenancy scenarios. Our results show that it effectively adapts to changing network conditions and fluctuating request rates at edge servers.
II Background
@@
TABLE I: Characteristics of common programmable accelerators.
@@
II-A Cloud and Edge Offloading
Computational offloading is a well-known approach where client devices that are resource-constrained offload some or all of their application workload to a remote server. In the case of mobile devices such as AR/VR headsets, tablets, or smartphones, offloading can use nearby edge resources to perform latency-sensitive tasks [4, 5]. While such edge offloading involves a network hop to the edge server (also known as the mobility penalty [7]), subsequent processing at the edge is assumed to be much faster since edge resources are much less constrained than the devices. In the case of IoT devices or ML applications, offloading to the cloud is also common where the abundant processing capacity of the cloud platform is utilized for processing the application workload [12]. For example, large language model (LLM) inference uses GPU clusters in the cloud to answer user queries, while IoT devices use cloud storage and processing to perform analytic and data processing tasks.
@@ II-A Cloud and Edge Offloading
II-B Hardware Accelerators
In recent years, a wide range of programmable accelerators has been developed to accelerate a range of compute tasks, particularly machine learning compute. Such accelerators have been deployed on client devices, edge, and cloud servers, enabling faster processing of accelerated tasks at each tier. Table 1 depicts the characteristics of several commonly deployed device and edge accelerators, ranging from low-end to high-end options. Specifically, the Intel Movidius Myriad X Vision Processing Unit (VPU) is designed to accelerate vision-based applications [13]. Google Edge Tensor Processing Unit (TPU) optimizes ML inference with the TensorFlow Lite framework on the edge [14]. NVIDIA TX2, Orin Nano, and A2 are all tailored for AI workloads such as deep learning and computer vision [15, 16, 17]. For these specific workloads, programmable accelerators can provide substantial performance improvements, delivering server-like capabilities for some compute-intensive tasks even on low-end devices. With their small form factor and energy-efficient design, integrating accelerators into mobile devices has become increasingly common. For example, Apple devices include a Neural accelerator to accelerate machine learning tasks locally on the device [10]. With the presence of accelerators, fast on-device processing of tasks such as machine learning inference has become feasible without the need to resort to edge offloading. Meanwhile, edge servers can also be equipped with GPUs, which serve as general-purpose accelerators for a range of tasks, or special-purpose accelerators for specific tasks. This enables offloaded tasks to also be accelerated at the edge, making edge offloading useful for more compute-intensive tasks that a device may not be able to handle using local resources.
II-C On-device Processing versus Edge Offloading
@@
Even with extensive offline profiling of execution times on both the device and the edge, selecting the optimal execution strategy remains challenging due to dynamic real-world conditions. On the one hand, mobile devices are subject to network variability due to fluctuating wireless signal strength, interference, and congestion in shared communication channels. Since edge offloading requires transmitting input data to an edge server, network fluctuations introduce unpredictability in network delays, affecting its overall performance. Additionally, mobile devices running applications with prolonged operating durations, such as live object detection on captured video, may experience battery constraints. When operating in low-power mode, the system may underclock its processing units to conserve energy, resulting in extended execution times. On the other hand, edge servers are multi-tenant systems shared by multiple devices, which can result in interference across users and higher latencies. Since devices may independently decide when and whether to offload tasks, the edge server’s load can fluctuate dynamically based on the number and nature of tasks being offloaded at any given time. Compared to traditional cloud data centers, edge servers are often more resource-constrained, limiting their ability to elastically scale applications under resource pressure. Workload spikes in multi-tenant scenarios can increase queuing delays, significantly increasing end-to-end execution times. Moreover, edge servers are commonly multiplexed across different applications to maximize edge resources utilization. At the same time, GPUs and other programmable accelerators designed for edge servers often lack isolation mechanisms such as GPU virtualization that are available in cloud-based GPUs [18], leading to performance interference between applications. As a result, execution times can become variable, making it challenging to determine whether to offload under dynamic workloads.
@@
III Model-driven Performance Comparison
In this section, we develop analytical queuing models for edge offloading and on-device processing, and derive closed-form expressions for their expected latencies to compare the two approaches and identify when one outperforms the other.
III-A Analytic Queuing Models
(a) Edge offloading queuing model
(b) On-device processing queuing model
Figure 1: Modeling request execution using (a) edge offloading and (b) on-device processing.
Our analytical models are derived from queuing theory, a well-established mathematical tool for modeling response times experienced by requests in computing systems. Queuing theory has previously been used to model the performance of edge and cloud applications [2, 19, 20], and will serve as the foundation for analyzing edge offloading and on-device processing. As discussed below, we model the end-to-end latency of a request—the duration of time from when the task enters the first queue until it exits the system—for both edge offloading and on-device processing using queuing theory.
Edge offloading. To model the end-to-end request execution time using edge offloading, we need to model the times spent by the request both at the device and at the edge server. We do so by modeling the device and the edge as two separate queuing systems as shown in Figure 1a. We assume that the request arrives at the device and is then forwarded (“offloaded”) to the edge server over the network. We need to model any network queuing delays and the network transmission time of the request from the device to the server. Upon arrival at the edge server, the request is queued for processing and is subsequently scheduled for execution. Since accelerators employ data parallelism to process requests, we model the accelerator at the edge server as a parallel system that processes each request with a degree of parallelism $k_{edge}$. Once the request processing completes, the result is sent back to the device, which incurs network queuing delay and transmission latency on the reverse path.
On-device processing. Similar to edge offloading, we model on-device request processing as a queuing system shown in Figure 1b. Incoming requests generated by the device first enter a local device queue and are then scheduled onto one of the accelerator cores of the device where they are processed with a degree of parallelism $k_{dev}$. Upon completion of processing, the result is returned to the local application and the request exits the system. This model captures an application where a client device generates tasks, utilizes its accelerator cores for local execution, and sends the result to the application.
@@ III-A Analytic Queuing Models
With the above notations, we now model the end-to-end latency of requests executed using each strategy:
$T_{edge}=w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$ (1)
@@
$T_{dev}=w_{dev}^{proc}+s_{dev}$ (2)
where $T_{edge}$ and $T_{dev}$ denote the end-to-end request latency under edge offloading and on-device processing, respectively.
III-B Deriving Service Times on Accelerators
Our analytic models require the service time of a workload as input to compute end-to-end response times. The service time reflects the computational demand on the target accelerator and indicates how quickly a request is processed on a given platform. It can be obtained through empirical profiling or a deep learning model that predicts workload execution times on accelerators. For empirical profiling, many applications already support function-level performance logging or provide profiling hooks [21], which can be used to derive the service time of a request by analyzing the corresponding function’s execution metrics. Alternatively, the workload can be profiled on the target hardware using benchmarking tools [22, 23]. In case of model-based prediction, a neural network can be trained to predict the service time of a model on a given target hardware. For example, prior work [24, 25, 26] estimates DNN inference latency by analyzing factors such as network structure and parameter size. For recurrent architectures such as Gated Recurrent Unit (GRU) networks and Long Short-Term Memory (LSTM) networks, existing latency estimation techniques [27, 28] incorporate factors such as temporal dependencies, hidden state dimensionality, and sequence unrolling behavior to accurately predict inference latency. While any of these prior techniques can be used with our analytic models, our experiments use a simple neural network model to predict the service times.
III-C Deriving and Analyzing End-to-End Latency
In this section we compute the expected latency for edge offloading and on-device processing. Our analysis focuses on applications that require processing results on the device such as AR/VR applications. Note that our models can also be extended to capture the performance of applications that require results on the edge. In the following, we leverage classical queuing theory to fundamentally understand the conditions under which one strategy outperforms the other.
III-C1 Comparing Edge Offload and On-Device Execution
We begin our analysis by examining the conditions under which edge offloading results in higher latency compared to on-device processing. For workloads that benefit from hardware acceleration such as those utilizing GPUs, processing times can be significantly reduced compared to processing on general-purpose CPUs. In the following discussion, we refer to such workloads as accelerator-driven workloads. While our analysis primarily focuses on DNN inference, the results generalize to other AI workloads as will be discussed in Section III-E. Table II summarizes the notation used in our models.
TABLE II: Notations used in analytic models.
@@ III-C1 Comparing Edge Offload and On-Device Execution
Lemma III.1.
For accelerator-driven workloads, edge offloading incurs a higher average end-to-end latency than on-device processing when {dmath} s_dev - s_edge ¡ λμdevnetdevnet- λ) + λμedgenetedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc )
@@
$w_{dev}^{proc}+s_{dev}<w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$ (3)
@@
$s_{dev}-s_{edge}<w_{dev}^{net}+w_{edge}^{net}+\frac{D_{req}}{B}+\frac{D_{res}}{B}+w_{edge}^{proc}-w_{dev}^{proc}$ (4)
As shown in [29], DNN inference on programmable accelerators often exhibits deterministic service times (denoted as $D$), as the number of operations per request remains constant. Since typical accelerators have a small degree of parallelism $k$ [30], the queuing system can be modeled as an $M/D/k/FCFS$ system, where request arrivals follow a Poisson process. While no closed-form solution exists for the expected wait time in such a system, a common approximation is to aggregate the service rate across the degree of parallelism $k$ [31, 32]. This reduces the model to an $M/D/1/FCFS$ system with aggregated service rate $k\mu$. Its expected wait time can be derived from the Pollaczek-Khinchin (P-K) formula [33, 19] and is given by
$E[w_{M/D/1}]=\frac{1}{2}\left(\frac{1}{k\mu-\lambda}-\frac{1}{k\mu}\right)$ (5)
@@
$E[w_{M/M/1}]=\frac{1}{\mu-\lambda}-\frac{1}{\mu}$ (6)
@@
Moreover, regardless of the execution strategy, the application sees the same arrival rate, i.e., $\lambda_{dev}^{proc}=\lambda_{edge}^{proc}=\lambda$. Substituting the above into (III-C1), we obtain {dmath} s_dev - s_edge ¡ λμdevnet(μdevnet- λ) + λμedgenet(μedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc ) which completes the proof. ∎
@@
Explanation. Service time, which is the time spent processing a request on an accelerator, is determined by the computational demand of the workload divided by the processing capacity of the hardware. For any given hardware configuration, requests with lighter demand have to be processed much faster at the edge to overcome the network penalty for edge offloading to be faster than on-device processing. For requests with lower computational demands, the difference in average service times $s_{dev}-s_{edge}$ shrinks proportionally due to the disparity in processing capacities between the device and the edge as the workload computational demand decreases. Further, since the service rate is the inverse of service times in queuing theory (i.e., $s=1/\mu$), the terms with $\mu_{dev}^{proc}$ and $\mu_{edge}^{proc}$ on the right hand side of the inequality in (III.1) will also become smaller since $\mu$ increases for lighter requests. The terms representing the network overheads remain unchanged since they depend on request size rather than the processing demand of the request. As a result, the inequality in (III.1) is more likely to hold true for requests with lighter processing demand. ∎
Remark III.2.
On-device processing is likely to outperform edge offloading on slower networks or when workloads have larger request or result payloads.
Explanation. The term $\frac{D_{req}+D_{res}}{B}$ in (III.1) increases when either the network bandwidth $B$ decreases or the request and result sizes $D_{req}$ and $D_{res}$ become larger, thereby increasing the right hand size of the inequality. In addition, longer transmission times reduce the service rates of the NIC at the device and edge ($\mu_{dev}^{net}$ and $\mu_{edge}^{net}$), further increasing the RHS. As a result, the inequality is more likely to hold for slower networks or tasks with higher network overheads, making on-device processing faster than edge offloading. ∎
Practical takeaways. Lemma III.1 establishes a bound on the average service time difference between device and edge for edge offloading to be more effective. The bound and the resulting inequality depends on several factors: (i) relative processing speeds of the device and edge, denoted by $s_{dev}$, $s_{edge}$ (and also the corresponding service rates $\mu_{dev}$ and $\mu_{edge}$), (ii) network speed, (iii) overall workload denoted by request rate $\lambda$, (iv) network transmission overheads. In general, for any specific workload, the edge processing capacity must be larger than the device processing capacity by a factor that overcomes the network penalty of offloading the workload; otherwise, on-device processing becomes preferable to offloading.